40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 203     Entry time: Wed Aug 28 11:04:00 2019
Author: Andy R. 
Type: Progress 
Category: BS BRDs 
Subject: BRD1 and 3 Updates 

BRD1 and BRD3 Resonance peaks after being in BS Suspension for ~1 month. Note that the BRD1 roll mode sagged a lot and the peaks were did not look sharp, so drift is probably unreliable. 

  BRD1 Roll BRD1 Bounce BRD3 Roll BRD3 Bounce
Measured f(Hz) 24.39 16.59 24.45 16.73
Drift in f (Hz) from 7/31 0.133 -0.10 0.075 0.133
Percent Drift 0.55% -0.62% 0.45% 0.55%

There appears to be less total drift when left in the BS suspension when compared to the BRDs that were measured more frequently. This seems to confirm the suspicion that excitation measurements cause some of the drift. 

ELOG V3.1.3-